# Quantum Hall and Light Responses in a 2D Topological Semimetal

### 06 Dec 2023 v2

@@ V Bulk-Edge correspondance
For the Kane-Mele model, as was shown in Fig. 1, the topological semimetal of interest in this paper can be understood as a superposition of two Haldane models associated to each spin polarization towards the axis of the in-plane magnetic field corresponding to the $r\mathbb{I}\otimes s_{x}$ term (or $\tilde{r}\mathbb{I}\otimes s_{x}$ term in the presence of interactions) in the Hamiltonian of Eq. (3). This magnetic field of small amplitude polarizes the system into two subsystems which are independent of each other as long as the $\mathbb{Z}_{2}$ symmetry of the Hamiltonian is preserved.
@@
The topological nature of this semimetal comes from the fact that each of the spin subsystem acts as if its chemical potential was shifted by ±$\pm r$ from its half-filling value, which leads for the bands to be either emptied or filled in a symmetric way between the two subsystems. This does not change however the topological nature of the bands themselves which keep their Berry curvature and their resulting Chern number. The semimetallic behavior thus does not come from a transition from a trivial to a topological insulator for which the topological nature of the band is not well defined. The topological nature of the bands is well defined here in terms of Chern number, but in the given context, the resulting band structure is gapless when taking the two spin populations together.
A way to probe this topological system is using spin-resolved ARPES measurement which should allow us to measure the two band structures of the upper panel of Fig. 4 separately and to resolve the bands lying inside the gap of each spin population, justifying hence the topological nature of this semimetal. However, the overall system will not develop a total quantized Hall conductance due to the presence of the bulk modes corresponding to the Fermi surface of each spin population, which, in a Landauer-Büttiker approach of the conductance, will result in an amplified value of the Hall conductance of the sample, not equal to the bulk value derived in Eq. (32). Hence, spin resolved ARPES measurements appear as the most simple and decisive way to probe the topological nature of this semimetal.
@@
Another way to probe the topological nature of this semimetal is using spin-polarized transport measurements at chemical potential $\mu$ which satisfies $|\mu\pm r|<|M-3\sqrt{3}t_{2}|$. Indeed, the spin polarization of the system in the $x$ direction is a well defined quantum number and in such a regime for the chemical potential, the mesoscopic Hall conductance of one spin population is quantized due to the unique presence of spin-polarized edge modes while the conductance of the other population is non quantized due to the presence of spin-polarized bulk modes at this regime of chemical potential. If we consider the graphic representations of the energy bands in Fig. 4, the system can thus be seen in this regime of chemical potential as a half-topological semimetal: one spin population is in the topological insulating phase while the other is in a metallic regime. We also emphasize here that the presence of a quantized conductance at the edges for one spin-population polarized along $x$ direction is also equivalent to a $\frac{1}{2}-\frac{1}{2}$ conductance for the two spin polarizations along $z$ direction which is precisely at the heart of the bulk-edge correspondence interpretation for the bilayer model [19, 27].
Similarly, spin resolved ARPES measurements on a generic spin Hall system should allow us to retrieve the band structure of each spin population independently, and thus to characterize the topological nature of each spin population on its own. This approach should allow us for exemple to exhibit the two band structures of the upper panel of Fig. 1 separately and to identify the topological nature of the spin Hall phase for the case of the Kane-Mele model. From Sec. III, for the topological semimetal, we also observe the equality
$\hat{C}_{\uparrow_{x},-}-\tilde{C}_{\downarrow_{x},+}=C_{\downarrow_{x},-},$ (50)
which is valid for all the ranges of applicability of Eq. (11) and formulates a correspondence between the integer $\mathbb{Z}$ topological invariant of the lowest-energy band and a $\mathbb{Z}_{2}$ spin topological number for the second-energy band.
This approach consisting in visualizing the band structure into distinct sectors of symmetry of the system brings powerful insight into the common definitions of the quantum Hall $\mathbb{Z}$ and of the quantum spin Hall $\mathbb{Z}_{2}$ invariants defined commonly in the literature. The general case would consist in a system made of several subsystems corresponding to distinct symmetry sectors of the symmetry of the system, and we can consider the system to be in a topological phase as long as one of the subsystems is itself in a topological phase. This tells us why the $\mathbb{Z}$ invariant is not relevant for the quantum spin Hall phase, because the two subsystems are of opposite Chern while still being separately in a topological phase, so that having the sum of their invariants being zero cannot help characterizing the topological nature of the phase. The $\mathbb{Z}_{2}$ invariant defined as the difference of the two invariants of the subsystems helps us finding a non zero topological invariant but this invariant is very much restricted to the precise context of subsystems being of opposite topology. Instead, at least on the theoretical level, one can consider more generally systems for which the symmetry sectors behave independently from each other and have their own topology, and neither the commonly defined $\mathbb{Z}$ or $\mathbb{Z}_{2}$ can help characterizing the topological phase entirely, as for exemple is the case of the following system
$\mathcal{H}_{{\bm{k}}}=\begin{pmatrix}d_{x}(\bm{k})\\ d_{y}(\bm{k})\\ M+\left(t_{2}+\tilde{t}_{2}\right)\sum_{j}\sin({\bf k}\cdot{\bf b}_{j})\end{pmatrix}\cdot\bm{\sigma}\otimes\mathbb{I}$ (54)
$+\left(t_{2}-\tilde{t}_{2}\right)\sum_{j}\sin({\bf k}\cdot{\bf b}_{j})\sigma_{z}\otimes s_{z}$ (55)
This model can realize different topological phases. This system would be in fact in a topological phase as long as $\left|\frac{M}{3\sqrt{3}t_{2}}\right|<1$ or $\left|\frac{M}{3\sqrt{3}\tilde{t}_{2}}\right|<1$. The case of the Kane-Mele model and the quantum spin Hall phase would correspond to $\tilde{t}_{2}=-t_{2}$ for which the $\mathbb{Z}_{2}$ spin topological number is relevant to characterize the topological phase, and the case $\tilde{t}_{2}=t_{2}$ would correspond to the usual bilayer Haldane model for which the $\mathbb{Z}$ number is relevant to characterize the topological phase. However, neither of these two invariants can help to characterize entirely the topological phase in the general context of $t_{2}\neq\pm\tilde{t}_{2}$. Hence, from general aspects, the topological phase of the system needs to be characterized for each symmetry sector independently, each having its own Chern number.
@@ VI Conclusion
To summarize, we have derived the quantum Hall conductivity of the recently identified two-dimensional topological nodal ring semimetal on the honeycomb lattice and obtained two distinct responses: a topological response of the lowest energy band which reveals an integer $\mathbb{Z}$ topological invariant and two Fermi surface terms from the two partially filled bands of opposite spins. We emphasize here that the Kubo formula is well applicable in the presence of a band-crossing effect as a result of orthogonality between spin quantum numbers of energy bands at the chemical potential. This orthogonality is also at the heart of the stability and protection of this phase of matter towards disorder and interaction effects which would only slightly dress the effect of the in-plane magnetic field. The $\mathbb{Z}$ topological invariant $C_{\downarrow_{x},-}$ associated to the lowest-energy band is also mathematically identical to a $\mathbb{Z}_{2}$ spin Chern number $\hat{C}_{\uparrow_{x},-}-\tilde{C}_{\downarrow_{x},+}$ for the partially filled bands. The specificity of the topological semimetal phase resides in the occurrence of gap closing region at the Fermi energy due to crossing between two bands which in the present case is also characterized through a $\mathbb{Z}_{2}$ spin Chern number. We have then evaluated the response to circularly polarized light, which is an equivalent marker of the topological response and of the quantum Hall conductivity for the Haldane model, and showed how this can reveal the topological response of the quantum Hall conductivity for the lowest band of the topological semimetal from the Dirac points.
This approach for the light-matter interaction resolved in the two Dirac valleys is also equivalent to a measure of the quantum Hall conductivity in the situation of a pair of one-half topological numbers related to the two-spheres’ model with a sphere associated to a $\uparrow_{z}$ or $\downarrow_{z}$ spin polarization. We have then justified that the formation of a half-topological semimetal can also be understood as one spin population along $x$ direction being in the topological insulating phase while the other is in a metallic regime resulting in a conductance at the edges quantized for one spin-polarization only. The quantized edge mode then has a 1/2 probability to be in the $\uparrow_{z}$, $\downarrow_{z}$ state associated to the original basis of the Hamiltonian. These results can be verified through spin-polarized transport at the edges and spin-polarized ARPES. These protected topological semimetals present then two different flows of charges, associated to a two-dimensional Fermi liquid in the bulk and a one-dimensional channel at the edge. It is also interesting to mention here that one-half topological numbers also find applications for the physics of interacting topological superconducting p-wave wires [43] and the physics of Majorana fermions [27, 44].
The bulk-edge correspondence yet hides some mysteries related to the classification of these robust, but at same time fragile, two-dimensional topological semimetals which may be addressed in future articles. Similarly, an interesting question may address the evolution of the present findings towards three dimensions.

### 27 Feb 2024 v3

@@ V Bulk-Edge correspondance
The topological semimetal of interest in this paper can also be understood as a superposition of two Haldane models associated here to each spin polarization towards the axis of the in-plane magnetic field corresponding to the $r\mathbb{I}\otimes s_{x}$ term in the Hamiltonian of Eq. (3). This magnetic field of small amplitude polarizes the system into two subsystems which are independent of each other as long as the $\mathbb{Z}_{2}$ symmetry of the Hamiltonian is preserved.
@@
The topological nature of this semimetal comes from the fact that each of the spin subsystem acts as if its chemical potential was shifted by ±$\pm r$ from its half-filling value, which leads for the bands to be either emptied or filled in a symmetric way between the two subsystems. This does not change however the topological nature of the bands themselves which keep their Berry curvature and their resulting Chern number. The topological nature of the bands is well defined here in terms of Chern number, but the resulting band structure is gapless when taking the two spin populations together. Spin resolved Angle Resolved PhotoEmission Spectroscopy (ARPES) appears as the most simple and decisive way to probe this semimetal. This should allow us to retrieve the band structure of each spin population independently, and thus to characterize the topological nature of each spin population on its own. This approach should also allow for example to reproduce the topological nature of the spin Hall phase for the case of the Kane-Mele model.
@@
Another way to probe this semimetal is using spin-polarized transport at chemical potential $\mu$ which satisfies $|\mu\pm r|<|M-3\sqrt{3}t_{2}|$. Indeed, the spin polarization of the system in the $x$ direction is a well defined quantum number and in such a regime for the chemical potential, the mesoscopic Hall conductance of one spin population is quantized due to the unique presence of spin-polarized edge modes while the conductance of the other population is non quantized due to the presence of spin-polarized bulk modes at this regime of chemical potential. If we consider the graphic representations of the energy bands in Fig. 2, the system can thus be seen in this regime of chemical potential as a half-topological semimetal: one spin population is in the topological insulating phase while the other is in a metallic regime. One quantized edge mode also agrees with the quantum Hall response of the lowest energy band. We also emphasize here that the presence of a quantized conductance at the edges for one spin-population polarized along $x$ direction is also equivalent to a $\frac{1}{2}-\frac{1}{2}$ conductance for the two spin polarizations along $z$ direction which is precisely at the heart of the bulk-edge correspondence interpretation for the bilayer model [22, 23] and the $\mathbb{Z}_{2}$ symmetry associated to the two spin polarizations along $z$ direction.
@@ VI Conclusion
To summarize, the quantum Hall conductivity of the recently identified two-dimensional topological nodal ring semimetal on the honeycomb lattice in one-layer graphene reveals a $\mathbb{Z}$ topological invariant $C_{\downarrow_{x},-}$ associated to the lowest-energy band and a $\mathbb{Z}_{2}$ spin Chern number $\hat{C}_{\uparrow_{x},-}-\tilde{C}_{\downarrow_{x},+}$ for the partially filled bands. The specificity of the topological semimetal phase resides in the occurrence of a gap closing region at the Fermi energy due to crossing between two bands which in the present case is also characterized through a $\mathbb{Z}_{2}$ spin Chern number in addition to the quantized quantum Hall response. We evaluated the response to circularly polarized light, which is an equivalent marker of the topological response and of the quantum Hall conductivity for the Haldane model, and showed how this can yet reveal the quantum Hall conductivity for the lowest-energy band of the topological semimetal from the frequency resonance corresponding to the energy gap at the Dirac points.
We have then justified that this system is a half-topological semimetal, i.e. one spin population along $x$ direction being in the topological insulating phase while the other is in a metallic regime resulting in a conductance at the edges quantized for one spin-polarization only. The quantized edge mode then has a 1/2 probability to be in the $\uparrow_{z}$, $\downarrow_{z}$ state associated to the original basis of the Hamiltonian. These results can be verified through spin-polarized transport at the edges and spin-polarized ARPES. The approach for the light-matter interaction resolved in the two Dirac valleys is also equivalent to a pair of one-half topological numbers associated to $\uparrow_{z}$ or $\downarrow_{z}$ spin polarizations, linking with the physical properties of the topological semimetal on the bilayer model [22, 23]. We also mention here that one-half topological numbers also find applications for the physics of interacting topological superconducting p-wave wires [37] and the physics of Majorana fermions [23, 38].
These protected topological semimetals present then two different flows of charges, associated to a two-dimensional Fermi liquid in the bulk and a one-dimensional (protected) channel(s) at the edges which may find applications in electronics and energy.